Circular No. 8429

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 V3 (SIDING SPRING) R. H. McNaught, Siding Spring Observatory, reports the discovery of

an object (discovery observation below) on images taken by G. J. Garradd with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey; McNaught later (Nov. 5 and 6) found the object to be slightly diffuse on CCD images taken with the Siding Spring 1.0-m f/8 reflector.

2004 UT  α_{2000}  δ_{2000}  Mag.
Nov. 3.39963  21^{h}45^{m}05.^{s}68  −45^{o}38′51′′.5  18.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-V33.

T = 2005 Jan. 22.836 TT  ω = 334.^{o}694
Ω = 354.211 }2000.0
q = 4.15415 AU  i = 51.393

COMET C/2004 V4 (NEAT) K. J. Lawrence, Jet Propulsion Laboratory, reports the NEAT dis-

covery of a comet (discovery observation below) on images taken with the 1.2-m Schmidt telescope at Palomar, the object showing very slight coma, but no tail. R. H. McNaught confirms that the object shows a slightly diffuse nuclear condensation and a 40′′-long fan tail centered in p.a. 280^{o} (the fan spanning 25^{o}) on 240-s images obtained with the Siding Spring 1-m reflector on Nov. 6.7.

2004 UT  α_{2000}  δ_{2000}  Mag.
Nov. 5.45377  4^{h}22^{m}36.^{s}96  +3^{o}39′28′′.3  18.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-V34.

T = 2005 Mar. 3.054 TT  ω = 42.^{o}667
Ω = 89.156 }2000.0
q = 1.27183 AU  i = 16.909

COMET 78P/GEHRELS Visual total magnitudes by A. Baransky, Pylypovychi, Ukraine (0.36-

m reflector): Sept. 18.93 UT, 12.0; Oct. 13.99, 11.0; Nov. 4.82, 10.4.

2004 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green